An Environmentalist is making a survey on the number of households that applies the principles of Waste Management in a certain community. Given the standard deviation of 26, how large a sample must be selected if we want to be 95% confident of finding whether the true mean differs from the sample mean by 1.5 hour?


1
Expert's answer
2022-06-09T14:49:49-0400

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c =1.96."

"E=z_c\times\dfrac{\sigma}{\sqrt{n}}""n\ge(\dfrac{z_c\times \sigma}{E})^2""n\ge(\dfrac{1.96\times 26}{1.5})^2"

"n=1155"